If sinx = 4/5 and siny= -12/13 , 0<x<Pi over 2, 3pi over 2 < y < 2pi, evaluate

a) cos (x+y)
b) tan (x-y)
1
Expert's answer
2020-11-01T17:15:03-0500


"\cos a = \pm \sqrt{1 - sin^2 a}, \tan a = \sin a / \cos a"


Precalculate:

"\cos x = \sqrt{1 - (4/5)^2} = 3/5, \tan x = 4/3"

"\cos y = \sqrt{1 - (-12/13)^2} = 5/13, \tan y = -12/5"


Answer:

"\begin{aligned}a) \cos(x+y) &= \cos x \cdot \cos y - \sin x \cdot \sin y \\
&= 3/5 \cdot 5/13 - 4/5 \cdot (-12/13) \\
&= 63/65\end{aligned}"


"\begin{aligned} b) \tan(x-y) &= (\tan x - \tan y) / (1 + \tan x \cdot \tan y) \\
&= (4/3 - (-12/5)) / (1 + 4/3 \cdot (-12/5)) \\
&= 56 / (15 - 48 ) = -56 / 33\end{aligned}"